# ENS-t-SNE: Embedding Neighborhoods Simultaneously t-SNE

### v2

@@ Abstract
[Uncaptioned image][Uncaptioned image][Uncaptioned image]
@@ 0.1 Motivation
Our proposed algorithm, named ENS-t-SNE, facilitates a challenging task within datasets: understanding the differences among individual data points, groups of points, or the dataset as a whole. In the typology of Brehmer and Munzner  [8], this is a “compare" task. There are two visualization comparison designs applicable to subspaces, as described by Gleicher et al. [16]: tjuxtaposition and superposition. Juxtaposition places points separately, as in a small multiple plot, but requires scanning to detect differences and similarities. Superpostion uses the same space to show two or more subspaces, resulting in more complex stimulus, which is closer to a direct comparison. ENS-t-SNE allows us access to both these comparison designs. The 3D embedding produced by ENS-t-SNE can be seen as a superpostion, by encoding each subspace from a different view of the object. ENS-t-SNE can provide juxtapostion with small multiple plots corresponding to projections for each subspace; see Fig. 4.
@@ 0.2 Our Contributions
We propose, describe and provide an implementation for ENS-t-SNE: a technique to perform dimension reduction on a high-dimensional dataset that captures multiple subspaces of interest in a single 3D embedding. Unlike the only prior work in this domain, which optimized global distance preservation (distances between all pairs of points) [19], we focus on preserving local relationships (clusters). We also visualize specific cluster relationships within the corresponding subspaces, with the help of 2D linear projections from the computed 3D embedding. Finally, we provide a seamless transition between the subspace views, which reduces the movement of the points from one view to the next, as the views are obtained by simple rotation of the 3D embedding; see Fig. 1. This is accomplished by generalizing the t-SNE algorithm [28] to simultaneously optimize multiple distance matrices, while also simultaneously optimizing projection views. In addition to the implementation, we provide a video (also in supplemental material).
@@ 1 Related Work
The computational complexity of t-SNE is high and speed improvements have been proposed [36]. Although the original paper proposes default values and ranges for the t-SNE hyperparameters (perplexity, learning rate, etc.), automatically selecting these parameters is also a topic of interest [9, 7]. A recent paper reviews t-SNE and applications thereof [15]. In our ENS-t-SNE algorithm we optimize a generalization of the cost function of t-SNE (eq. 1) per projection, as detailed in section 2.
@@
where $\sigma^{2}$ is the MDS stress function. MPSE aims to preserve global distances: all pairwise distances in the input high-dimensional space must match all pairwise distances in low dimensional space. In contrast, ENS-t-SNE aims to preserve in low dimensional space only distances between nearby points in high dimensional space.
@@ 2.1 ENS-T-SNE Algorithm
Figure 3: MPSE applied to the synthetic data described in section 3. Recall that we should see clusters based on texture, shape, and color in the three views. MPSE fails in capturing this information by mixing clusters in the color and shape views. In general, clusters are much more well seperated for the same data in Fig 2.
@@ 3 Experiments on Synthetic Data
Figure 4: The Palmer’s Penguins dataset captured by MPSE. Note how MPSE mixes the blue and orange clusters.
@@
Experiments: Since the goal of ENS-t-SNE is to preserve local structures, this dataset suits its purpose and serves as a good example to show experimentally how the proposed algorithm works. The aim is to show that ENS-t-SNE preserves the local structures of the dataset, that is, each of the perspectives recovers the corresponding original clusters.
We run ENS-t-SNE for datasets generated according to Section 3. For the first experiment, we fix the number of perspectives to be $M=3$, the number of datapoints $N=400$, the number of clusters per perspective $NC_{m}=2$ for $1\leq m\leq 3$. We create the distance matrices for each perspective by formula (13) described in Section 3 with $d_{in}=1$ and $d_{out}=2$ and $\sigma^{2}=0.1$. We run the ENS-t-SNE algorithm, summarized in Section 2.1, for the distance matrices and a fixed perplexity value of $40$. The result are demonstrated in Figure 1.
@@ 4.1 Palmer’s Penguins
Applying standard t-SNE on this dataset produces six distinct clusters, one for each sex-species pair; see Fig. 6(a). However, we can obtain a more fine-grained information by utilizing ENS-t-SNE and two subspaces: the 4-dimensional physical attributes and the 1-dimensional sex attribute. Specifically, we create two pairwise distance matrices, the first one based on the 4D physical attributes which have numerical values. The second one is based on penguin’s sex, where the distance between same sex penguins is 0 and the distance between different sex penguins is 1. Running ENS-t-SNE with these distance matrices produces the embedding in Fig. 4(a).
The first view (middle) aims to capture the physical attributes and the second view (right) aims to capture the sex information. While ENS-t-SNE also clusters the data into 6 distinct clusters in 3D, the positions of these clusters are more meaningful. In the first view (Fig. 4b) the algorithm has captured physical attributes which are largely correlated with the 3 penguin species colored blue for Adelie, orange for Gentoo, and green for Chinstrap. Meanwhile, the second view (Fig. 4c) shows two large clusters based on sex: circles for female penguins and rectangles for male penguins.
While one could run t-SNE on each subdimension separately and plot them as small multiples, such a visualization does not capture the correspondence between datapoints in the two independent embeddings. In the ENS-t-SNE embedding, each point belongs to two clusters; one for its species and one for its sex. In an interactive environment, one can follow a datapoint from one projection to the other. In other words, there is a transition between the two views in three dimensions that is missing when using small multiples. Comparing the t-SNE and ENS-t-SNE embeddings of the dataset we can see that unlike in t-SNE, the positions of the clusters in ENS-t-SNE are meaningful. For example, the two orange clusters in t-SNE embedding in Figure 6 are far from each other, while they are close in the ENS-t-SNE embedding in Figure 4. Thus both datapoint positions and cluster positions are more meaningful in the ENS-t-SNE embedding.
@@ 4.2 Food Composition Dataset
Figure 7: ENS-t-SNE applied to the USDA Food Composition dataset. (a): The 3D embedding found by ENS-t-SNE, where each of the three classes have been seperated. (b): The first projection corresponding to the water+lipids subspace. The blue and orange clusters (grains and vegetables, fruits and drinks) have been separated, with several blue points mixed in both the orange and red clusters. (c): The second projection corresponding to the proteins+B vitamins subspace. Here the red cluster (meats) has been well separated but the blue and orange cluster contain similar protein and B vitamin components so have been mixed into two clusters, the smaller of which corresponds to dairy and the larger to other meatless foods.
@@ 5 Quantitative Evaluation
For each metric, we compute the 3D values with respect to all distances, as well as the per-perspective values with respect to each subspace. For the following definitions, let $N$ be the size of a dataset (number of points), and $K$ be a parameter for the size of a neighborhood. For trustworthiness, continuity, and neighborhood hit we use $K=7$ for all evaluations, as in [12].
@@
where $Ne_{i}^{K}$ is the $K$ nearest neighborhood of point $i$ in the embedding, and $l_{i}$ is the label of point $i$.
We note that the food composition dataset does not have ground truth labels, but there is good evidence that our clustering is accurate. The clustering has a high silhouette score and is visually validated in Fig 6 and Fig. 7.
@@
where $X_{i}$ is the embedded position of point $i$ and $d_{i,j}$ is the distance between points $i,j$ in the high dimensional space. Low stress indicates good distance preservation. Note that MPSE directly optimizes stress for each view, so we expect it to often outperform ENS-t-SNE.
Table 1 shows the results. Note how ENS-t-SNE tends to outperform MPSE on the continuity, trustworthiness, and NH scores for the 2d views indicated that as expected ENS-t-SNE is more reliably capturing local structures in the projected views. Low stress is not the goal of this algorithm, so having high stress compared to MPSE is acceptable.
@@
Stability. Although we cannot directly quantitatively compare to t-SNE (or other single perspective dimension reduction embeddings), we can measure how similar a set of projections are. It is known that for many comparative tasks it is desirable to have as little change as possible while still being faithful to the data. This notion is often referred to as stability or preservation of the mental map [3, 11].
@@ 6 Limitations
While MPSE [19] focuses on simultaneously capturing global distances between objects and ENS-t-SNE aims to capture local neighborhoods, other approaches for dimension reduction, such as UMAP [29], optimize both at the same time. It would be worthwhile to quantitatively verify the extend to which these goals can be realized by the different approaches.
The utility of ENS-t-SNE depends on finding interesting subspaces/subdimensions and combinations thereof. We have not yet considered automating the process by using approaches such as those for subspace clustering as part of the ENS-t-SNE pipeline. Setting up the distance matrices that ENS-t-SNE needs can be done in different ways, as illustrated in the different examples in the paper. Evaluating different approaches and automating the process remains to be done.
While we expect that 3D ENS-t-SNE embeddings might be easier to interpret and work with, compared to small multiple type visualizations (with or without linked views), we are not aware of human-subject studies to validate this intuition.
@@ 7 Conclusions and Future Work
An interesting direction that we began to explore is to extend the objective function such that each perspective shows the t-SNE embedding for different values of perplexities; see the supplemental material.
Another possible application is using ENS-t-SNE to visualize image datasets, based on different parts of the input images. We include some preliminary results for the MNIST dataset in the supplemental material.
@@ Appendix C Visualizing the Effect of Perplexity
Figure 16: Synthetic data generated for the quantitative experiment. The two clusters in the first 10 dimensions are seperated using color, the next 10 dimensions are denoted by shape, and the final 10 denoted by fill.
@@
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" alt="[LOGO]">

### v3

@@ 0.1 Motivation
Our proposed algorithm, named ENS-t-SNE, facilitates a challenging task within datasets: understanding the differences among individual data points, groups of points, or the dataset as a whole. In the typology of Brehmer and Munzner  [8], this is a “compare" task. There are two visualization comparison designs applicable to subspaces, as described by Gleicher et al. [16]: juxtaposition and superposition. Juxtaposition places points separately, as in a small multiple plot, but requires scanning to detect differences and similarities. Superposition uses the same space to show two or more subspaces, resulting in more complex stimulus, which is closer to a direct comparison. ENS-t-SNE allows us access to both these comparison designs. The 3D embedding produced by ENS-t-SNE can be seen as a superposition, by encoding each subspace from a different view of the object. ENS-t-SNE can provide juxtaposition with small multiple plots corresponding to projections for each subspace, replacing standard independent projections of subspaces; see Fig. ENS-t-SNE: Embedding Neighborhoods Simultaneously t-SNE.
@@ 0.2 Our Contributions
We propose, describe and provide an implementation for ENS-t-SNE: a technique to perform dimension reduction on a high-dimensional dataset that captures multiple subspaces of interest in a single 3D embedding. Unlike the only prior work in this domain, which optimized global distance preservation (distances between all pairs of points) [19], we focus on preserving local relationships (clusters). We also visualize specific cluster relationships within the corresponding subspaces, with the help of 2D linear projections from the computed 3D embedding. Finally, we provide a seamless transition between the subspace views, which reduces the movement of the points from one view to the next, as the views are obtained by simple rotation of the 3D embedding; see Fig. 1. This is accomplished by generalizing the t-SNE algorithm [28] to simultaneously optimize multiple distance matrices, while also simultaneously optimizing projection views. In addition to the implementation, we provide a publicly available web demonstration  [^3] https://jacoblmiller.github.io/ENS-t-SNE-web/. We remark that, unlike t-SNE, where the relative positions of clusters are usually meaningless, ENS-t-SNE does better in preserving these positions. It embeds clusters in 3D in a manner that ensures the corresponding 2D projections contain the corresponding clusters, compelling the 3D clusters to have meaningful positions. We experimentally verify this and observe it visually for several datasets.
@@ 1 Related Work
The computational complexity of t-SNE is high and speed improvements have been proposed [36]. Although the original paper proposes default values and ranges for the t-SNE hyperparameters (perplexity, learning rate, etc.), automatically selecting these parameters is also a topic of interest [9, 7]. recent paper reviews t-SNE and applications thereof [15]. In our ENS-t-SNE algorithm we optimize a generalization of the cost function of t-SNE (eq. 1) per projection, as detailed in section 2.
@@
where $\sigma^{2}$ is the MDS stress function. MPSE aims to preserve global distances: all pairwise distances in the input high-dimensional space must match all pairwise distances in low-dimensional space. In contrast, ENS-t-SNE aims to preserve in low dimensional space only local neighborhoods.
@@ 2.1 ENS-T-SNE Algorithm
Figure 3: MPSE applied to the synthetic data from Section 3. Recall that we should see clusters based on texture, shape, and color in the three views. MPSE fails in capturing this information by mixing clusters in the color and shape views. In general, these clusters are better separated in Fig 2.
@@ 3 Experiments on Synthetic Data
Figure 4: The Palmer’s Penguins dataset captured by MPSE. (a): The full 3D embedding. (b): The projection capturing physical characteristics, encoded by color. (c): The embedding capturing penguin sex, encoded by shape. MPSE mixes the blue and orange clusters also squared and circled clusters.
Refer to captionRefer to caption
@@
Experiments: Since the goal of ENS-t-SNE is to preserve local structures, this dataset suits its purpose and serves as a good example to show experimentally how the proposed algorithm works. The aim is to show that ENS-t-SNE preserves the local structures of the dataset, that is, each of the perspectives recovers the corresponding original clusters. We conduct experiments on two synthetic datasets: the first with $400$ datapoints, three perspectives, and two clusters each; the second with $1000$ datapoints, three perspectives, and two, three, and four clusters.
We run ENS-t-SNE for datasets generated according to Section 3. For the first experiment, we fix the number of perspectives to be $M=3$, the number of datapoints $N=400$, the number of clusters per perspective $NC_{m}=2$ for $1\leq m\leq 3$. We create the distance matrices for each perspective by formula (13) described in Section 3 with $d_{in}=1$ and $d_{out}=2$ and $\sigma^{2}=0.1$. We run the ENS-t-SNE algorithm, summarized in Section 2.1, for the distance matrices and a fixed perplexity value of $40$. The result is demonstrated in Figure 1.
@@ 4.1 Palmer’s Penguins
Applying standard t-SNE on this dataset produces six distinct clusters, one for each sex-species pair; see Fig. 6(a). However, we can obtain a more fine-grained information by utilizing ENS-t-SNE and two subspaces: the 4-dimensional physical attributes and the 1-dimensional sex attribute. Specifically, we create two pairwise distance matrices, the first one based on the 4D physical attributes which have numerical values. The second one is based on penguin’s sex, where the distance between same sex penguins is 0 and the distance between different sex penguins is 1. Running ENS-t-SNE with these distance matrices produces the embedding in Fig. ENS-t-SNE: Embedding Neighborhoods Simultaneously t-SNE(a). The first view (middle) aims to capture the physical attributes and the second view (right) aims to capture the sex information. While ENS-t-SNE also clusters the data into 6 distinct clusters in 3D, the positions of these clusters are more meaningful. In the first view (Fig. ENS-t-SNE: Embedding Neighborhoods Simultaneously t-SNEb) the algorithm has captured physical attributes which are largely correlated with the 3 penguin species colored blue for Adelie, orange for Gentoo, and green for Chinstrap. Meanwhile, the second view (Fig. ENS-t-SNE: Embedding Neighborhoods Simultaneously t-SNEc) shows two large clusters based on sex: circles for female penguins and rectangles for male penguins.
While one could run t-SNE on each subdimension separately and plot them as small multiples, such a visualization does not capture the correspondence between datapoints in the two independent embeddings. In the ENS-t-SNE embedding, each point belongs to two clusters; one for its species and one for its sex. In an interactive environment, one can follow a datapoint from one projection to the other. In other words, there is a transition between the two views in three dimensions that is missing when using small multiples. Comparing the t-SNE and ENS-t-SNE embeddings of the dataset we can see that unlike in t-SNE, the positions of the clusters in ENS-t-SNE are meaningful. For example, the two orange clusters in t-SNE embedding in Figure 6 are far from each other, while they are close in the ENS-t-SNE embedding in Figure ENS-t-SNE: Embedding Neighborhoods Simultaneously t-SNE. Thus both datapoint positions and cluster positions are more meaningful in the ENS-t-SNE embedding.
In Fig. 4, we compare MPSE embedding of Palmer’s Penguins dataset to ENS-t-SNE (Fig. ENS-t-SNE: Embedding Neighborhoods Simultaneously t-SNE) using the same variables. In the first view (Fig. 4(b)), blue and orange points are mixed, and in the second view (Fig. 4(c)), squared and circled shapes are mixed. ENS-t-SNE, however, offers superior visualization with clear separation of components in each view. In Fig. 5, we demonstrate the 3D embeddings of the same dataset using MDS, t-SNE, and UMAP. While these methods are not directly comparable to ENS-t-SNE and MPSE, the intent is to highlight the benefits of simultaneous embedding when features are somewhat separated. In all embeddings, points roughly form six clusters, but there is no distinct separation by both color and shape.
@@ 4.2 Food Composition Dataset
Figure 7: ENS-t-SNE applied to the USDA Food Composition dataset. (a): The 3D embedding found by ENS-t-SNE, where each of the three classes have been separated. (b): The first projection corresponding to the water+lipids subspace. The blue and orange clusters (grains and vegetables, fruits and drinks) have been separated, with several blue points mixed in both the orange and red clusters. (c): The second projection corresponding to the proteins+B vitamins subspace. Here the red cluster (meats) has been well separated but the blue and orange cluster contain similar protein and B vitamin components so have been mixed into two clusters, the smaller of which corresponds to dairy and the larger to other meatless foods.
@@ 5 Quantitative Evaluation
For each metric, we compute the 3D values with respect to all distances, as well as the per-perspective values with respect to each subspace. For the following definitions, let $N$ be the size of a dataset (number of points), and $K$ be a parameter for the size of a neighborhood. For trustworthiness, continuity, and neighborhood hit we use $K=7$ for all evaluations, as in [12]. We use the same perplexity value for each dataset as described in their figures: perplexity of 40 for the Palmer’s Penguins and value of 30 for the Auto-MPG and Food Composition dataset. These values were chosen by investigating a range in (20, 100) and selecting the best quantitative embedding.
@@
where $Ne_{i}^{K}$ is the $K$ nearest neighborhood of point $i$ in the embedding, and $l_{i}$ is the label of point $i$. We note that the food composition dataset does not have ground truth labels, but there is good evidence that our clustering is accurate. The clustering has a high silhouette score and is visually validated in Fig 6 and Fig. 7.
@@
where $X_{i}$ is the embedded position of point $i$ and $d_{i,j}$ is the distance between points $i,j$ in the high dimensional space. Low stress indicates good distance preservation. Note that MPSE directly optimizes stress for each view, so we expect it to often outperform ENS-t-SNE. Table 1 shows the results of these metrics, averaged over 10 trials each. Note how ENS-t-SNE tends to outperform MPSE on the continuity, trustworthiness, and NH scores for the 2d views indicating that as expected ENS-t-SNE is more reliably capturing local structures in the projected views. Low stress is not the goal of this algorithm, so having high stress compared to MPSE is acceptable. ENS-t-SNE can achieve lower stress in one view than MPSE, since MPSE minimizes the average stress per projection, meaning that ENS-t-SNE may produce unbalanced projections with respect to stress where MPSE will produce more balanced projections.
@@
Stability. Although we cannot directly quantitatively compare to t-SNE (or other similar dimension reduction embeddings), we can measure how similar a set of projections are. It is known that for many comparative tasks it is desirable to have as little change as possible while still being faithful to the data. This notion is often referred to as stability or preservation of the mental map [3, 11]. We show that ENS-t-SNE is more stable when computing a set of projections of subspaces than standard t-SNE on the same set of subspaces.
@@ 5.1 Scalability
Refer to captionRefer to caption
Refer to captionRefer to caption
@@ 6 Limitations
While MPSE [19] focuses on simultaneously capturing global distances between objects and ENS-t-SNE aims to capture local neighborhoods, other approaches for dimension reduction, such as UMAP [29], optimize both at the same time. It would be worthwhile to quantitatively verify the extend to which these goals can be realized by the different approaches. The utility of ENS-t-SNE depends on finding interesting subspaces/subdimensions and combinations thereof. We have not yet considered automating the process by using approaches such as those for subspace clustering as part of the ENS-t-SNE pipeline. Setting up the distance matrices that ENS-t-SNE needs can be done in different ways, as illustrated in the different examples in the paper. Evaluating different approaches and automating the process remains to be done.
While we expect that 3D ENS-t-SNE embeddings might be easier to interpret and work with, compared to small multiple type visualizations (with or without linked views), we are not aware of human-subject studies to validate this intuition. Although we were inspired by 3D physicalizations one can “walk around" and interact with, a thorough human subjects study is required to verify that this method supports tasks effectively.
@@ 7 Conclusions and Future Work
An interesting direction that we began to explore is to extend the objective function such that each perspective shows the t-SNE embedding for different values of perplexities; see the supplemental material. Another possible application is using ENS-t-SNE to visualize image datasets, based on different parts of the input images. We include some preliminary results for the MNIST dataset in the supplemental material.
@@ Appendix B Visualizing MNIST Dataset by ENS-t-SNE
Refer to captionRefer to captionRefer to captionRefer to captionRefer to captionRefer to captionRefer to captionRefer to captionRefer to captionRefer to caption
@@
Refer to captionRefer to captionRefer to captionRefer to captionRefer to captionRefer to caption
@@ Appendix C Visualizing the Effect of Perplexity
Refer to captionRefer to captionRefer to captionRefer to captionRefer to caption
@@
Figure 16: Synthetic data generated for the quantitative experiment. The two clusters in the first 10 dimensions are separated using color, the next 10 dimensions are denoted by shape, and the final 10 denoted by fill.
Refer to captionRefer to captionRefer to caption